AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 20 of 33  ELOG logo
ID Date Author Subject
  281   Fri Feb 25 09:58:08 2022 GHC, CB, ML, EMCurrent status

10:58
Started new run: 77

Nozzle clogged more when trying warming again.
Warming up, venting top and bottom to clean nozzle.
 

  280   Fri Feb 25 08:14:47 2022 George, EstherCurrent status

9:14
Beam intensity around ~7.5e7

Pressures OK
System check OK
White rabbit fail diff: 24
Write speed: ~1.8Mb/s
Leakage current 3.49uA
Cryostat Setpoint: 121.5K

  279   Fri Feb 25 06:57:09 2022 George, EstherCurrent status

08:00
Beam seems to be back
Intensity ~7.8e7

Pressures OK
System check OK
White rabbit fail diff: 24
Write speed: ~1.3Mb/s
Leakage current 3.49uA
Cryostat Setpoint: 121.5K

Attachment 1: 2022-02-25_08-03-21leakage.png
2022-02-25_08-03-21leakage.png
  278   Fri Feb 25 06:28:12 2022 George, EstherCurrent Status

07:30
Some signals on the beam monitor, seems they are optimising it.

System Check OK
Temps OK
White Rabbit fail diff: 24
Leakage Current: 3.48uA
Pressures OK
Data rate ~1Mb/s

Attachment 1: 2022-02-25_07-33-44exRate.png
2022-02-25_07-33-44exRate.png
Attachment 2: 2022-02-25_07-35-11xyTotal.png
2022-02-25_07-35-11xyTotal.png
  277   Fri Feb 25 05:04:16 2022 Rui-Jiu Chen, Alexcheck

Still  no beam in CRING.

Motor fully out position.

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 217 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

Attachment 1: Screenshot_from_2022-02-25_06-05-12.png
Screenshot_from_2022-02-25_06-05-12.png
Attachment 2: Screenshot_from_2022-02-25_06-05-32.png
Screenshot_from_2022-02-25_06-05-32.png
Attachment 3: Screenshot_from_2022-02-25_06-06-12.png
Screenshot_from_2022-02-25_06-06-12.png
Attachment 4: Screenshot_from_2022-02-25_06-06-49.png
Screenshot_from_2022-02-25_06-06-49.png
Attachment 5: Screenshot_from_2022-02-25_06-07-24.png
Screenshot_from_2022-02-25_06-07-24.png
Attachment 6: Screenshot_from_2022-02-25_06-07-52.png
Screenshot_from_2022-02-25_06-07-52.png
Attachment 7: Screenshot_from_2022-02-25_06-08-15.png
Screenshot_from_2022-02-25_06-08-15.png
Attachment 8: Screenshot_from_2022-02-25_06-09-09.png
Screenshot_from_2022-02-25_06-09-09.png
Attachment 9: Screenshot_from_2022-02-25_06-09-50.png
Screenshot_from_2022-02-25_06-09-50.png
Attachment 10: Screenshot_from_2022-02-25_06-10-35.png
Screenshot_from_2022-02-25_06-10-35.png
Attachment 11: Screenshot_from_2022-02-25_06-11-27.png
Screenshot_from_2022-02-25_06-11-27.png
Attachment 12: Screenshot_from_2022-02-25_06-12-09.png
Screenshot_from_2022-02-25_06-12-09.png
Attachment 13: Screenshot_from_2022-02-25_06-12-41.png
Screenshot_from_2022-02-25_06-12-41.png
Attachment 14: Screenshot_from_2022-02-25_06-13-23.png
Screenshot_from_2022-02-25_06-13-23.png
  276   Fri Feb 25 03:20:29 2022 Rui-Jiu Chen, AlexCurrent status

Gleb Vorobyev tried for few hours to fix the ion source but without success.

So, there will be no beam in CRYRING during this night shift. Today at about 08:00 the efforts to fix the ion source will be restarted.

  275   Fri Feb 25 03:01:11 2022 Rui-Jiu Chen, Alexcheck

Still  no beam in CRING.

Motor fully out position.

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 236 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

Attachment 1: Screenshot_from_2022-02-25_04-02-04.png
Screenshot_from_2022-02-25_04-02-04.png
Attachment 2: Screenshot_from_2022-02-25_04-02-17.png
Screenshot_from_2022-02-25_04-02-17.png
Attachment 3: Screenshot_from_2022-02-25_04-03-14.png
Screenshot_from_2022-02-25_04-03-14.png
Attachment 4: Screenshot_from_2022-02-25_04-03-52.png
Screenshot_from_2022-02-25_04-03-52.png
Attachment 5: Screenshot_from_2022-02-25_04-04-21.png
Screenshot_from_2022-02-25_04-04-21.png
Attachment 6: Screenshot_from_2022-02-25_04-05-04.png
Screenshot_from_2022-02-25_04-05-04.png
Attachment 7: Screenshot_from_2022-02-25_04-05-35.png
Screenshot_from_2022-02-25_04-05-35.png
Attachment 8: Screenshot_from_2022-02-25_04-06-25.png
Screenshot_from_2022-02-25_04-06-25.png
Attachment 9: Screenshot_from_2022-02-25_04-07-06.png
Screenshot_from_2022-02-25_04-07-06.png
Attachment 10: Screenshot_from_2022-02-25_04-07-48.png
Screenshot_from_2022-02-25_04-07-48.png
Attachment 11: Screenshot_from_2022-02-25_04-10-05.png
Screenshot_from_2022-02-25_04-10-05.png
Attachment 12: Screenshot_from_2022-02-25_04-11-13.png
Screenshot_from_2022-02-25_04-11-13.png
Attachment 13: Screenshot_from_2022-02-25_04-12-21.png
Screenshot_from_2022-02-25_04-12-21.png
Attachment 14: Screenshot_from_2022-02-25_04-13-16.png
Screenshot_from_2022-02-25_04-13-16.png
  274   Fri Feb 25 01:12:49 2022 Rui-Jiu Chen, Alex Check

Still  no beam in CRING.

Motor fully out position.

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 243 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

Attachment 1: Screenshot_from_2022-02-25_02-13-46.png
Screenshot_from_2022-02-25_02-13-46.png
Attachment 2: Screenshot_from_2022-02-25_02-14-56.png
Screenshot_from_2022-02-25_02-14-56.png
Attachment 3: Screenshot_from_2022-02-25_02-16-05.png
Screenshot_from_2022-02-25_02-16-05.png
Attachment 4: Screenshot_from_2022-02-25_02-16-54.png
Screenshot_from_2022-02-25_02-16-54.png
Attachment 5: Screenshot_from_2022-02-25_02-17-29.png
Screenshot_from_2022-02-25_02-17-29.png
Attachment 6: Screenshot_from_2022-02-25_02-17-39.png
Screenshot_from_2022-02-25_02-17-39.png
Attachment 7: Screenshot_from_2022-02-25_02-17-57.png
Screenshot_from_2022-02-25_02-17-57.png
Attachment 8: Screenshot_from_2022-02-25_02-19-05.png
Screenshot_from_2022-02-25_02-19-05.png
Attachment 9: Screenshot_from_2022-02-25_02-19-17.png
Screenshot_from_2022-02-25_02-19-17.png
Attachment 10: Screenshot_from_2022-02-25_02-20-01.png
Screenshot_from_2022-02-25_02-20-01.png
Attachment 11: Screenshot_from_2022-02-25_02-21-36.png
Screenshot_from_2022-02-25_02-21-36.png
Attachment 12: Screenshot_from_2022-02-25_02-22-17.png
Screenshot_from_2022-02-25_02-22-17.png
Attachment 13: Screenshot_from_2022-02-25_02-23-40.png
Screenshot_from_2022-02-25_02-23-40.png
  273   Fri Feb 25 00:00:40 2022 Rui-Jiu Chen, AlexCheck

There are some problems with ion source.  No beam in CRYING.

Motor fully out position.

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 230 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

Attachment 1: Screenshot_from_2022-02-25_01-02-48.png
Screenshot_from_2022-02-25_01-02-48.png
Attachment 2: Screenshot_from_2022-02-25_01-02-58.png
Screenshot_from_2022-02-25_01-02-58.png
Attachment 3: Screenshot_from_2022-02-25_01-04-09.png
Screenshot_from_2022-02-25_01-04-09.png
Attachment 4: Screenshot_from_2022-02-25_01-06-45.png
Screenshot_from_2022-02-25_01-06-45.png
Attachment 5: Screenshot_from_2022-02-25_01-07-30.png
Screenshot_from_2022-02-25_01-07-30.png
Attachment 6: Screenshot_from_2022-02-25_01-07-55.png
Screenshot_from_2022-02-25_01-07-55.png
Attachment 7: Screenshot_from_2022-02-25_01-08-22.png
Screenshot_from_2022-02-25_01-08-22.png
Attachment 8: Screenshot_from_2022-02-25_01-09-24.png
Screenshot_from_2022-02-25_01-09-24.png
Attachment 9: Screenshot_from_2022-02-25_01-10-03.png
Screenshot_from_2022-02-25_01-10-03.png
Attachment 10: Screenshot_from_2022-02-25_01-12-00.png
Screenshot_from_2022-02-25_01-12-00.png
Attachment 11: Screenshot_from_2022-02-25_01-15-34.png
Screenshot_from_2022-02-25_01-15-34.png
Attachment 12: Screenshot_from_2022-02-25_01-16-22.png
Screenshot_from_2022-02-25_01-16-22.png
Attachment 13: Screenshot_from_2022-02-25_01-17-24.png
Screenshot_from_2022-02-25_01-17-24.png
Attachment 14: Screenshot_from_2022-02-25_01-17-58.png
Screenshot_from_2022-02-25_01-17-58.png
  272   Thu Feb 24 23:15:25 2022 Rui-Jiu Chen, AlexProblems with ion souce.
  271   Thu Feb 24 22:45:08 2022 Jordan, George, Rui-Jiu Chencheck

- Motor fully out position.

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 236 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

 

Attachment 1: Screenshot_from_2022-02-24_23-49-31.png
Screenshot_from_2022-02-24_23-49-31.png
Attachment 2: Screenshot_from_2022-02-24_23-54-01.png
Screenshot_from_2022-02-24_23-54-01.png
Attachment 3: Screenshot_from_2022-02-24_23-54-57.png
Screenshot_from_2022-02-24_23-54-57.png
Attachment 4: Screenshot_from_2022-02-24_23-56-46.png
Screenshot_from_2022-02-24_23-56-46.png
Attachment 5: Screenshot_from_2022-02-24_23-58-18.png
Screenshot_from_2022-02-24_23-58-18.png
Attachment 6: Screenshot_from_2022-02-24_23-59-10.png
Screenshot_from_2022-02-24_23-59-10.png
Attachment 7: Screenshot_from_2022-02-24_23-59-32.png
Screenshot_from_2022-02-24_23-59-32.png
Attachment 8: Screenshot_from_2022-02-25_00-00-48.png
Screenshot_from_2022-02-25_00-00-48.png
Attachment 9: Screenshot_from_2022-02-25_00-02-54.png
Screenshot_from_2022-02-25_00-02-54.png
Attachment 10: Screenshot_from_2022-02-25_00-04-35.png
Screenshot_from_2022-02-25_00-04-35.png
Attachment 11: Screenshot_from_2022-02-25_00-09-26.png
Screenshot_from_2022-02-25_00-09-26.png
Attachment 12: Screenshot_from_2022-02-25_00-11-10.png
Screenshot_from_2022-02-25_00-11-10.png
Attachment 13: Screenshot_from_2022-02-25_00-13-30.png
Screenshot_from_2022-02-25_00-13-30.png
Attachment 14: Screenshot_from_2022-02-25_00-14-11.png
Screenshot_from_2022-02-25_00-14-11.png
  270   Thu Feb 24 14:31:25 2022 JM, ML, CNBeam Tuning Run 76

15:30

1.5 MeV/u beam currently being tuned

Target is ON, setpoint = 135 K

System checks OK - white rabbit fail diff ~ 9

Temperatures OK

Merger OK

Data rate ~ 1400 kb/sec

Leakage current ~ 3.41 uA

Aida 1 and 2 missing channels from previous days appear working on stat layout 2 and xy pixel total plot (attachment 3)

17:12

Beam tuning complete. See attached Ex rate histogram.

The cryo compressor is being turned OFF for around 1 hour to remove any frozen nitrogen which may be the cause of the increased pressures observed.

Beam OFF to do so.

System checks OK - white rabbit fail diff ~ 9

Temperatures OK

Merger OK

Data rate ~ 200 kb/sec

Compressor OFF, setpoint on controller is 273 K

18:51

Pressure in the ring recovered quickly after closing valves to gas jet (attachment 5)

Increasing temperature caused a large spike in pressure in the inlets. E1 turbo briefly shut down due to the pressure increase, since restarted and pressure back in the -4 mbar range. (attachment 6)

Temperature being decreased

Beam back ON

data rate ~ 1500 kb/sec

merger OK

Temperatures OK

System checks OK - white rabbit diff ~ 12

21:58

After cooling the jet target, expected density is not achieved. E4 pressure is ~4E-10 mbar compared to ~2E-8mbar previously. Temperature reduced to 127K E4 pressure increases however stilll within -10 mbar range.

With beam on target we see rutherford peak at ~3 MeV, however it is much reduced.

Beam OFF to investigate inside the cave.

Decreasing temperature of nozzle in 0.5 k steps to see effect on E4 and S1 dump.

 

Attachment 1: 2022-02-24_15-35-53-Ex-rate-histogram.png
2022-02-24_15-35-53-Ex-rate-histogram.png
Attachment 2: 2022-02-24_15-39-08-leakage-current.png
2022-02-24_15-39-08-leakage-current.png
Attachment 3: 2022-02-24_15-44-06-aida1-5-stat.png
2022-02-24_15-44-06-aida1-5-stat.png
Attachment 4: 2022-02-24_17-07-52-xyTotal.png
2022-02-24_17-07-52-xyTotal.png
Attachment 5: vacplot24.png
vacplot24.png
Attachment 6: gasjet-vacplot.png
gasjet-vacplot.png
  269   Thu Feb 24 09:43:16 2022 JG, MScheck

Beam is still beeing tuned. Target is off.

 

- Motor on manual

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 947 kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide checks OK

- Spectra ok

SW5

4 plots uploaded

 

Attachment 1: Screenshot_from_2022-02-24_10-42-50.png
Screenshot_from_2022-02-24_10-42-50.png
Attachment 2: Screenshot_from_2022-02-24_10-42-29.png
Screenshot_from_2022-02-24_10-42-29.png
Attachment 3: Screenshot_from_2022-02-24_10-42-13.png
Screenshot_from_2022-02-24_10-42-13.png
Attachment 4: Screenshot_from_2022-02-24_10-41-45.png
Screenshot_from_2022-02-24_10-41-45.png
Attachment 5: Screenshot_from_2022-02-24_10-41-27.png
Screenshot_from_2022-02-24_10-41-27.png
Attachment 6: Screenshot_from_2022-02-24_10-41-04.png
Screenshot_from_2022-02-24_10-41-04.png
Attachment 7: Screenshot_from_2022-02-24_10-40-53.png
Screenshot_from_2022-02-24_10-40-53.png
Attachment 8: Screenshot_from_2022-02-24_10-40-42.png
Screenshot_from_2022-02-24_10-40-42.png
Attachment 9: Screenshot_from_2022-02-24_10-40-26.png
Screenshot_from_2022-02-24_10-40-26.png
Attachment 10: Screenshot_from_2022-02-24_10-40-02.png
Screenshot_from_2022-02-24_10-40-02.png
Attachment 11: Screenshot_from_2022-02-24_10-39-37.png
Screenshot_from_2022-02-24_10-39-37.png
Attachment 12: Screenshot_from_2022-02-24_10-39-22.png
Screenshot_from_2022-02-24_10-39-22.png
Attachment 13: Screenshot_from_2022-02-24_10-39-10.png
Screenshot_from_2022-02-24_10-39-10.png
  268   Thu Feb 24 08:30:36 2022 JG, MSAIDA5 calibration test failed - solved

UPDATE: THE PROBLEM IS SOLVED FOR NOW:

Doing the calib for the 5th time eventually worked.

------------

The ADC check showed that the calib test for aida5 failed.

We tried to do the calibration as described in the fail message, but it seems to not work. See attachement.

Attachment 1: Screenshot_from_2022-02-24_09-09-36.png
Screenshot_from_2022-02-24_09-09-36.png
Attachment 2: Screenshot_from_2022-02-24_09-29-24.png
Screenshot_from_2022-02-24_09-29-24.png
  267   Thu Feb 24 08:15:23 2022 JG, MScheck

Beam is still beeing tuned. Target is off.

There is a problem with ADC calibration of aida5: see entry https://elog.ph.ed.ac.uk/CARME/268

 

- Motor on manual

SW3

- Merger OK

- Screenshot uploaded

- Data rate : ??? kbytes/s

SW4

- current plot uploaded

SW2

- T OK

- system wide check OK (ADC calibration test for AIDA5 failed, see above)

- Statistics ok

- Spectra ok

SW5

4 plots uploaded

 

Attachment 1: Screenshot_from_2022-02-24_09-16-09.png
Screenshot_from_2022-02-24_09-16-09.png
Attachment 2: Screenshot_from_2022-02-24_09-14-51.png
Screenshot_from_2022-02-24_09-14-51.png
Attachment 3: Screenshot_from_2022-02-24_09-14-31.png
Screenshot_from_2022-02-24_09-14-31.png
Attachment 4: Screenshot_from_2022-02-24_09-14-10.png
Screenshot_from_2022-02-24_09-14-10.png
Attachment 5: Screenshot_from_2022-02-24_09-13-51.png
Screenshot_from_2022-02-24_09-13-51.png
Attachment 6: Screenshot_from_2022-02-24_09-10-41.png
Screenshot_from_2022-02-24_09-10-41.png
Attachment 7: Screenshot_from_2022-02-24_09-10-25.png
Screenshot_from_2022-02-24_09-10-25.png
Attachment 8: Screenshot_from_2022-02-24_09-10-08.png
Screenshot_from_2022-02-24_09-10-08.png
Attachment 9: Screenshot_from_2022-02-24_09-09-36.png
Screenshot_from_2022-02-24_09-09-36.png
Attachment 10: Screenshot_from_2022-02-24_09-26-41.png
Screenshot_from_2022-02-24_09-26-41.png
Attachment 11: Screenshot_from_2022-02-24_09-26-00.png
Screenshot_from_2022-02-24_09-26-00.png
Attachment 12: Screenshot_from_2022-02-24_09-25-11.png
Screenshot_from_2022-02-24_09-25-11.png
Attachment 13: Screenshot_from_2022-02-24_09-24-52.png
Screenshot_from_2022-02-24_09-24-52.png
  266   Thu Feb 24 07:25:53 2022 OH (Remote)Recover of system after crash
08:00 Woke up to message from Carlo explaining the server had crashed.
      Anydesked into the machine.
      Merger was showing just data transfer it was not showing Transfer => Merger

      Went through the steps detailed in https://elog.ph.ed.ac.uk/DESPEC/36 to recover it.
      After following these steps was again able to forward data from the merger.
      Tested and it now sends on data again.

      Setup the screen session for the caen bias and restarted the logger program.
      Restarted the gnuplot script to plot leakage current
      Started the online monitoring software and browser.

      Have posted in the chat that the DAQ is now running again.
  265   Thu Feb 24 05:09:47 2022 CB, PE, EMInlet valve at target closed + problems with turbopump

The turbo pump E1 is continuing to draw too much current and the rotation frequency has steadily decreased from 817 Hz at the beginning of the night shift to 800 Hz now. The frequency recovered after the temperature was increased.

Both of the valves (inlet and dump) have been closed to improve the vacuum in the ring.

  264   Thu Feb 24 04:17:35 2022 CB, TD (remotely)MIDAS DAQ no longer working

DAQ server crashed around 24:00

Rebooted server but could not could not rescue remotely - could not contact raspberry pi in control of FEE bias.

Entered cave, rebooted raspberry pi. Would not accept remote connections.

Had to manually start DHCP service on carme-gsi using

systemctl start dhcpd

 

Raspi could then be connected.

Started DAQ, TapeServer and Merger as per usual. Tapeserver refuses writing to disk.

Checked merged - very odd timestamp situation. See attached. Timestamp look OK in the GSI WR tab. See attached.

Power cycled server and FEES

Timestamp issue solved. All are now reading 0x16 etc.

 

However tapeserver still does not write to disk. Reasons unclear. Possibly Merger is not merging and not passing data on?

Have to stop data collection until solved.

 

Attachment 1: Bad_Time_2022-02-24_04-34-12.png
Bad_Time_2022-02-24_04-34-12.png
Attachment 2: BadTime_WRTime_2022-02-24_04-34-53.png
BadTime_WRTime_2022-02-24_04-34-53.png
  263   Wed Feb 23 22:29:50 2022 PE, EM, CB (remotely)check - 23:30

Motor position ok

Data rate: 1030kbytes/s

T ok

System wide check ok

plots uploaded

 

DAQ crashed soon afterwards

  262   Wed Feb 23 21:27:29 2022 TD, CNcheck

Motor position ok

Data rate: 1030kbytes/s

T ok

System wide check ok

plots uploaded

 

Attachment 1: Screenshot_from_2022-02-23_22-29-36.png
Screenshot_from_2022-02-23_22-29-36.png
Attachment 2: Screenshot_from_2022-02-23_22-30-24.png
Screenshot_from_2022-02-23_22-30-24.png
Attachment 3: Screenshot_from_2022-02-23_22-33-23.png
Screenshot_from_2022-02-23_22-33-23.png
Attachment 4: Screenshot_from_2022-02-23_22-36-57.png
Screenshot_from_2022-02-23_22-36-57.png
Attachment 5: Screenshot_from_2022-02-23_22-37-22.png
Screenshot_from_2022-02-23_22-37-22.png
Attachment 6: Screenshot_from_2022-02-23_22-39-50.png
Screenshot_from_2022-02-23_22-39-50.png
Attachment 7: Screenshot_from_2022-02-23_22-41-08.png
Screenshot_from_2022-02-23_22-41-08.png
Attachment 8: Screenshot_from_2022-02-23_22-42-51.png
Screenshot_from_2022-02-23_22-42-51.png
Attachment 9: Screenshot_from_2022-02-23_22-43-16.png
Screenshot_from_2022-02-23_22-43-16.png
Attachment 10: Screenshot_from_2022-02-23_22-43-39.png
Screenshot_from_2022-02-23_22-43-39.png
Attachment 11: Screenshot_from_2022-02-23_22-46-52.png
Screenshot_from_2022-02-23_22-46-52.png
ELOG V3.1.4-unknown